OpenStudy (anonymous):

The intensity I of light bulb varies inversely as the square of the distance d from the light bulb. Suppose I is 250 w/m^2 when the distance is 5 m. How much farther would it be to a point where the intensity is 40 w/m^2?

OpenStudy (anonymous):

since they vary inversely.. I * d = constant so in first case I1*d1 = 250*5 = 1250 so I2 * d2 should also be 1250 in second case 40 * d2 = 1250 d2 = 1250/40 = 31.25 d1 = 5 d2 = 31.25 So it would be 31.25 - 5 = 26.25 m further to the point where the intensity is 40 w/m^2